# Document Valuation in LLM Summaries: A Cluster Shapley Approach

### v2

@@ 2 Related Literature
Our research relates to and contributes to multiple streams of work, including game theory, LLMs, and marketing. First, our work relates to the growing literature on LLM-based summarization. Recent research on summarization has shown that summaries produced by LLMs like GPT-4 achieve comparable or superior factual accuracy, coherence, and overall quality compared to human annotators in news summarization tasks (pu2023summarization). Indeed, text summarization has evolved from traditional extractive methods like TextRank (mihalcea2004textrank) to more sophisticated abstractive approaches powered by LLMs. RAG frameworks (lewis2020retrieval), which combine information retrieval with LLM generation, have shown significant improvements in factual accuracy and information currency (gao2023retrieval; jiang-etal-2023-active). Recent developments include GraphRAG (edge2024local), which enhances retrieval performance using graph-based representations. While summarization technologies continue to evolve rapidly, our work addresses the fundamental challenge of document valuation that persists across summarization approaches. The document valuation framework we develop in §6.2 is designed to be agnostic to specific summarization techniques, ensuring its applicability even as LLM and RAG technologies advance.
@@
Next, our work builds on the game-theoretic concept of Shapley values (shapley1953value), which has gained significant traction in machine learning for quantifying feature importance (lundberg2017unified) and data valuation (jia2019towards; ghorbani2019data). [^5] Apart from the Shapley value approach, alternative approaches for calculating the value of contributors in cooperative tasks have also been proposed. Some of the well-known approaches include the leave-one-out approach (cook1977detection) and the influence function approach (barshan2020relatif; han2020explaining; guo2020fastif). However, these methods, neglecting high-order interactions, may result in undesirable data valuation. For example, when there are two or more similar contributors, and tend to assign nearly zero value to similar documents. For example, consider a scenario where we have three documents, A, B, and C, where A and B have the same high-quality content and C contributes nothing to the overall value. Intuitively, we should expect A and B to have the same valuation and C to have zero value. However, the leave-one-out approach will assign zero value to both A and B – if we leave A out and calculate the value generated by B and C, the value would almost be the same as that with A (and vice-versa for B). Thus, if we believe this approach and assume that both A and B have no incremental value and drop them both from the set of contributors, the overall value generated by the set will go to zero! In sum, these approaches are unable to correctly account for the marginal contributions of each player. This is the reason why Shapley has emerged as the dominant paradigm for valuing contributors in cooperative settings. While prior studies have applied Shapley values primarily to supervised learning tasks and feature attribution, our work represents, to our knowledge, the first application of Shapley values to document valuation in LLM-based summarization systems. Our work also relates to the work on approximation algorithms for computing Shapley values since it is well-known that computing exact Shapley values is computationally challenging (mann1960values; ghorbani2019data). However, these earlier work treats each observation as independent and do not leverage potential similarities between contributors. In contrast, we explicitly account for the semantic similarity between documents in our proposed Cluster Shapley algorithm, which allows us to achieve high accuracy at relatively low computation costs. We expect that our general idea of using a clustering approach to capture the similarity between documents/observations in Shapley calculations can be extended beyond LLM summarization settings and applied more broadly to other cases where similarity measures between observations exist.
@@
While the notion of computing Shapley values for groups of data points has been previously mentioned – e.g., ghorbani2019data briefly discuss group-level Shapley values in their Section 3.2, and corder2019shapley explore a hierarchical clustering approach combined with exact Shapley calculations—we make three distinct contributions. First, we explicitly introduce an approximate method for computing cluster-level Shapley values, offering practical gains in computational cost. Second, we formally analyze the trade-off between approximation error and computational complexity, whereas corder2019shapley only provide theoretical results under restrictive assumptions (e.g., K-means with two clusters). Third, to our knowledge, we are the first to use LLM-generated text embeddings to guide the clustering, which better captures semantic relationships and is especially suited to our task.
@@
More broadly, recent works such as wang2024economic have used the notion of Shapley value to quantify the value of large corpora to pretraining by approximating Shapley values through repeated fine-tuning (an expensive and often infeasible process). In contrast, our method targets post-training applications – specifically, LLM-driven summarization tasks, such as AI search or QA, where the model is fixed and attribution occurs at inference time. Because our approach approximates Shapley values based on LLM-generated embeddings, it avoids retraining and scales more efficiently with the number of input documents. This distinction is crucial: wang2024economic deals with corpus-level attribution during LLM training; we focus on attribution over documents presented as context to an already-trained LLM, making our method more practical for interactive, high-frequency applications.
Finally, our work relates to the literature on news aggregators in marketing and economics. News aggregators function similarly to AI-based summarization, since users may consume content directly on the aggregator’s platform without visiting the publishers’ websites. A stream of research has examined whether this substitutive effect reduces traffic to publishers or whether aggregators can also serve as a discovery channel that increases exposure. For instance, mayzlin2012link find that blogs can build an audience by effectively promoting/linking to rival sites under certain conditions. dellarocas2013media develop a game-theoretic model and show that content aggregators can both benefit and harm content creators and consumers. jeon2016news investigate competition among online newspapers in the presence of aggregators. Empirically, both calzada2020news and athey2021impact find that news aggregators can exert a market-expansion effect, increasing visits to news outlets, especially for smaller publishers. Further, recent work by amaldoss2023can explores how publishers might collaborate and compete with news aggregators, and song2023does empirically examine the effects of carrying news on user engagement with non-news content on social media. Given that AI-based summarization is functionally similar to content aggregators, our Cluster Shapley approach for valuing source documents can also be used by content aggregators and newspapers/content websites to formalize revenue-sharing arrangements.
@@ 3 Problem Definition
We define the problem from the perspective of a platform that has access to $D$ original documents generated by different producers. We do not make any distributional assumptions on $D$. These documents are not necessarily i.i.d. and may contain overlapping information and vary in the quantity and quality of content. Users arrive at the platform and query for some information from the platform using queries $q$, drawn from a distribution $g(q)$. The platform generates a response to each query $q$ based on the $D$ documents using an LLM-based summarization model $A(q,D)$. We can view A(q,D){A(q,D)} as a black box that takes a dataset $D$ of any size between 0 and $\infty$ to generate a summary in response to query $q$. Note that in practice, the platform may choose to only use a subset of documents ($S_{q}\subseteq D$) that are most relevant to the query for the summarization process; that is, we allow for cases where all documents are not relevant to all queries. In such cases, the summarization process is denoted by $A(q,S_{q})$.
The quality or performance of a summarization is denoted by $v(q,A(q,S_{q}))$. Intuitively, this score captures the extent to which the user finds the summary useful or valuable. The performance score $v$ can be treated as a black-box oracle that takes the query and summary as input and returns a score. In practice, $v(q,A(q,S_{q})$ can be obtained in a multitude of ways. It could be actual scores collected from user surveys on how helpful they find a given summary to be (e.g., rating of helpfulness, fraction of upvotes). Alternatively, it could be helpfulness scores based on an LLM model, where an independent LLM agent does the scoring instead of human agents. This can be a viable option in settings where collecting user responses is costly and/or slow; indeed, recent research has shown that LLM ratings tend to align with user ratings in many situations (kang2023llms). It could also represent implicit helpfulness scores based on user behavior, which are commonly used in the information retrieval and search literature to measure the relevance of a given document/link, e.g., whether the user clicked on the summary, the time spent reading the summary (liu_2009; yoganarasimhan_2020).
@@
We define $\phi_{i}(q)$ as the value of document $i$ for a given query $q$. This value function depends on the subset of documents retrieved for query $q$ (i.e., $S_{q}\subseteq D$), the summary function of documents $A(q,S)$, and the resulting performance score function $v(q,A(q,S))$. Formally, $\phi_{i}(q)=\phi_{i}(q,S_{q},A,v).$ This formulation has two key advantages. First, documents that are irrelevant to a given query—i.e., not included in $S_{q}$—naturally receive zero value: $\phi_{i}(q)=0$ if $i\notin S_{q}$. Second, this setup allows the platform to aggregate document values across queries while respecting both the relative importance of documents and the frequency of the queries they support. For instance, a document that is essential for answering a rare query may be less valuable overall than a document that provides moderate value across many common queries.
Our goal is to develop a document valuation approach that satisfies two properties:
@@
    Summarization Procedure Agnostic: The approach should be agnostic to the specifics of the summarization process, $A(\cdot)$, used by the platform. That is, it should generalize across RAG systems. While we present a standard RAG implementation in $\S$6.2, numerous alternatives exist—from simpler methods to more sophisticated frameworks like TextRank, GraphRAG, and DRAG (mihalcea2004textrank; edge2024local; zhang2025lexical). As LLM technologies rapidly evolve, our document valuation framework is designed to remain effective, regardless of advancements in summarization, ensuring broad applicability across current and future implementations.
@@
    Evaluation Process Agnostic: The approach should apply to any evaluation method ($v(\cdot)$). As discussed earlier, many explicit and implicit approaches for scoring summaries exist. Different business use cases may have access to different evaluation approaches. For example, search engines (e.g., Perplexity or Google) usually only have implicit feedback/evaluation, whereas question-answering websites or review aggregators may have more explicit feedback on the helpfulness of reviews. We would like our algorithm to be agnostic to the exact approach used. In our empirical context, we use a prompt-based LLM approach for evaluation; see $\S$6.3.
In sum, our goal is to develop a solution concept that is agnostic to the details of the generative summarization model used ($A(\cdot)$) and the evaluation procedure ($v(\cdot)$) and is broadly applicable across a variety of domains and business applications of LLM summaries, e.g., revenue attribution as discussed in $\S$4.2.
4 Solution Concept: Shapley Framework for Document Valuation
This section presents a principled framework for addressing the problem of document valuation in LLM summaries, as introduced in $\S$3. In $\S$4.1, we introduce Shapley value, a game-theory concept for assigning value to individual documents in cooperative settings, and explain its relevance to our context. In $\S$4.2, we discuss the business application of using Shapley value for revenue attribution. While theoretically appealing, computing exact Shapley values is not practically applicable in most settings (including ours). To address this, $\S$4.3 introduces an efficient approximation algorithm—Cluster Shapley—which leverages document similarity measures to reduce the computational burden. Finally, in $\S$4.4, we analyze the theoretical properties of this algorithm and provide guarantees on approximation error and computation complexity.
@@ 4.1 Shapley Value
We now introduce the Shapley value formula along with a concise, self-contained explanation of the framework. Based on $\S$3, recall that our goal is to find a document valuation function $\rho_{i}\left(D,A(\cdot),v(\cdot),g(\cdot)\right)\in\mathbb{R}$ to quantify the value of document $i$ in set $D$. To obtain this valuation, we need to first accurately estimate the query-level document valuation function $\phi_{i}\left(D,q,A(q,S_{q}),v(q,A(q,S_{q})\right)$. Henceforth, we denote this value function as $\phi_{i}(q)$ because the retrieval process $S_{q}$, the LLM-based summarization process $A(q,S_{q})$, and the performance score function $v(q,A(q,S_{q}))$ are all uniquely defined by $q$.
Following the standard Shapley literature, we now adapt and present the four properties of $\phi_{i}(q)$ within the context of our LLM-summarization problem to guarantee equitable document valuation:
@@
    If the efficiency property is not enforced, the value function $\phi$ is only determined up to a proportional constant (see Theorem 2.1 in (ghorbani2019data)).
@@
    Note that the combination of two queries $q_{1}+q_{2}$ does not imply that the two queries are merged into a single new query. Instead, it represents a setting where there are two distinct queries being processed (this definition naturally extends to any finite number of queries, not just two). For example, consider two queries: one on quality ($q_{1}$) and another on price ($q_{2}$). The value of document $i$ under the quality query is denoted by $\phi_{i}(q_{1})$, calculated using the performance score function $v(q_{1},A(q_{1},S))$. Similarly, $\phi_{i}(q_{2})$ represents the value of document $i$ under the price query, based on the performance score function $v(q_{2},A(q_{2},S))$. The linearity property asserts that if we sum the values obtained from these separate queries, i.e., $\phi_{i}(q_{1})+\phi_{i}(q_{2})$, the result is equivalent to the value of document $i$ calculated under a new, combined performance score function $v(q_{1},A(q_{1},S))+v(q_{2},A(q_{2},S))$. This combined value is $\phi_{i}(q_{1}+q_{2})$.
While other desirable properties are worth discussing, these four – Efficiency, Symmetry, Null Document, and Linearity – uniquely determine the document value function $\phi(q)$, Shapley value (shapley1953value); and no additional conditions are necessary. [^6] These four properties (axioms) are independent of any specific summarization $(A)$ or evaluation $(v)$, which are laid out in our problem definition. These properties originate from Shapley’s work on cooperative game theory (shapley1953value). This is a foundational result in cooperative game theory. We refer interested readers to Shapley’s seminal paper for a formal proof, and henceforth only focus on how these properties apply to our context and their importance in our setting.
Under the above four properties, the Shapley value $\phi_{i}(q)$ for a document $i\in S_{q}\subseteq D$ is uniquely expressed as the expected marginal contribution of document $i$ across all possible coalitions:
@@
Recall that for $i\notin S_{q}$, the Shapley value of $i$ is $\phi_{i}(q)=0$ because only documents in $S_{q}$ are used for summarization. Documents outside $S_{q}$ have no contribution to the query and thus receive a Shapley value of zero. The Shapley value formula can also be written based on the original document set $D$ as:
@@
However, many of these evaluation score calculations are redundant since any document outside $S_{q}$ does not increase the performance score. In fact, this formula defined on $D$ is equivalent to the formula defined only on $S_{q}$ in Equation (4). The equivalence follows directly from the permutation-based definition in the earlier paragraph, as all permutations of $D\setminus S_{q}$ do not affect the performance score.
A natural concern in attribution systems is whether document providers can strategically manipulate their content to inflate their assigned value. This issue is not new: even in traditional search engines like Google, publishers have long engaged in “mimicking the winner” strategies—creating content that resembles highly ranked pages to improve visibility. While such behavior is theoretically possible in RAG-based systems, it is even more difficult to execute in practice in these systems. Modern LLM-powered platforms (e.g., Perplexity, ChatGPT, Google’s AI Overview) do not disclose their specific RAG systems to external document providers. Moreover, these systems do not present explicit ranked lists but instead synthesize information into summaries, making it harder for publishers to infer their position or influence outcomes. Retrieval in such systems is governed by proprietary, black-box algorithms that incorporate a range of dynamic signals, including semantic embeddings, recency, diversity, and user engagement. This level of opacity is a deliberate design choice, observed not only in LLM systems but also in recommender platforms and digital advertising ecosystems, where hiding algorithmic details is a known strategy to deter gaming (wang2023algorithmic).

### v3

@@ 2 Related Literature
Our research relates to and contributes to multiple streams of work, including game theory, LLMs, and business applications of AI systems. First, our work relates to the growing literature on LLM-based summarization. Recent research on summarization has shown that summaries produced by LLMs like GPT-4 achieve comparable or superior factual accuracy, coherence, and overall quality compared to human annotators in news summarization tasks (52). Indeed, text summarization has evolved from traditional extractive methods like TextRank (45) to more sophisticated abstractive approaches powered by LLMs. RAG frameworks (36), which combine information retrieval with LLM generation, have shown significant improvements in factual accuracy and information currency (19; 31). Recent developments include GraphRAG (14), which enhances retrieval performance using graph-based representations. While summarization technologies continue to evolve rapidly, our work addresses the fundamental challenge of document valuation that persists across summarization approaches. The document valuation framework we develop is designed to be agnostic to specific summarization techniques, ensuring its applicability even as LLM and RAG technologies advance.
@@
Next, our work builds on the game-theoretic concept of Shapley values (62), which has gained significant traction in machine learning for quantifying feature importance (40) and data valuation (30; 20). [^6] Apart from the Shapley value approach, alternative approaches for calculating the value of contributors in cooperative tasks have also been proposed. Some of the well-known approaches include the leave-one-out approach (10) and the influence function approach (3; 27; 25). However, these methods, neglecting high-order interactions, may result in undesirable data valuation. For example, when there are two or more similar contributors, and tend to assign nearly zero value to similar documents. For example, consider a scenario where we have three documents, A, B, and C, where A and B have the same high-quality content and C contributes nothing to the overall value. Intuitively, we should expect A and B to have the same valuation and C to have zero value. However, the leave-one-out approach will assign zero value to both A and B – if we leave A out and calculate the value generated by B and C, the value would almost be the same as that with A (and vice-versa for B). Thus, if we believe this approach and assume that both A and B have no incremental value and drop them both from the set of contributors, the overall value generated by the set will go to zero! In sum, these approaches are unable to correctly account for the marginal contributions of each player. This is the reason why Shapley has emerged as the dominant paradigm for valuing contributors in cooperative settings. While prior studies have applied Shapley values primarily to supervised learning tasks and feature attribution, our work represents, to our knowledge, the first application of Shapley values to document valuation in LLM-based summarization systems. Our work also relates to the work on approximation algorithms for computing Shapley values since it is well-known that computing exact Shapley values is computationally challenging. Without additional structural assumptions on the value function beyond bounded range, any unbiased estimator must rely on sampling permutations or coalitions; simple Monte Carlo (random-permutation) estimators are unbiased and achieve concentration rates that match information-theoretic lower bounds up to constants (5; 41). Specifically, with value range bounded by $[a,b]$, Hoeffding-style arguments yield a sample complexity of $O\!\left(\frac{(b-a)^{2}}{\varepsilon^{2}}\log\frac{1}{\delta}\right)$ to obtain an $\varepsilon$-accurate estimate with probability $1-\delta$, and Maleki et al. (41) give matching lower bounds. Alternative speedups such as linear regression surrogates, kernel/coalitional weighting, or grouping assumptions can reduce computation but introduce modeling assumptions that may bias attribution when violated (18; 2; 20).
@@
In our proposed Cluster Shapley algorithm, we explicitly account for the semantic similarity between documents, which allows us to achieve high accuracy at relatively low computation costs. We expect that our general idea of using a clustering approach to capture the similarity between documents/observations in Shapley calculations can be extended beyond LLM summarization settings and applied more broadly to other cases where similarity measures between observations exist. While the notion of computing Shapley values for groups of data points has been previously mentioned – e.g., Ghorbani and Zou (20) briefly discuss group-level Shapley values in their Section 3.2, and Corder and Decker (11) explore a hierarchical clustering approach combined with exact Shapley calculations-we make two contributions. First, we formally analyze the trade-off between approximation error and computational complexity, whereas Corder and Decker (11) only provide theoretical results under restrictive assumptions (e.g., K-means with two clusters). Second, to our knowledge, we are the first to use LLM-generated text embeddings to guide the clustering, which better captures semantic relationships and is especially suited to our task.
@@
More broadly, recent works such as Wang et al. (67) have used the notion of Shapley value to quantify the value of large corpora to pretraining by approximating Shapley values through repeated fine-tuning (an expensive and often infeasible process). In contrast, our method targets post-training applications – specifically, LLM-driven summarization tasks, such as AI search or QA, where the model is fixed and attribution occurs at inference time. Because our approach approximates Shapley values based on LLM-generated embeddings, it avoids retraining and scales more efficiently with the number of input documents. This distinction is crucial: Wang et al. (67) deals with corpus-level attribution during LLM training; we focus on attribution over documents presented as context to an already-trained LLM, making our method more practical for interactive, high-frequency applications.
@@ 3 Problem Definition
We define the problem from the perspective of a platform that has access to $D$ original documents generated by different producers. We do not make any distributional assumptions on $D$. These documents are not necessarily i.i.d. and may contain overlapping information and vary in the quantity and quality of content. Users arrive at the platform and query for some information from the platform using queries $q$, drawn from a distribution $g(q)$. The platform generates a response to each query $q$ based on the $D$ documents using an LLM-based summarization model $A(q,D)$. We can view A(q,D){A(q,D)} as a black box that takes a dataset $D$ of any size from 0 to $\infty$ to generate a summary in response to query $q$. Note that in practice, the platform may choose to only use a subset of documents ($S_{q}\subseteq D$) that are most relevant to the query for the summarization process; that is, we allow for cases where all documents are not relevant to all queries. In such cases, the summarization process is denoted by $A(q,S_{q})$.
The quality or performance of a summarization is denoted by $v(q,A(q,S_{q}))$. Intuitively, this score captures the extent to which the user finds the summary useful or valuable. The performance score $v$ can be treated as a black-box that takes the query and summary as input and returns a score. In practice, $v(q,A(q,S_{q}))$ can be obtained in a multitude of ways. It could be actual scores collected from user surveys on how helpful they find a given summary to be (e.g., rating of helpfulness, fraction of upvotes). Alternatively, it could be helpfulness scores based on an LLM model, where an independent LLM agent does the scoring instead of human agents. This can be a viable option in settings where collecting user responses is costly and/or slow; indeed, recent research has shown that LLM ratings tend to align with user ratings in many situations (32; 8). It could also represent implicit helpfulness scores based on user behavior, which are commonly used in the information retrieval and search literature to measure the relevance of a given document/link, e.g., whether the user clicked on the summary, the time spent reading the summary (38; 71).
@@
We define $\phi_{i}(q,S_{q},A,v)$ as the value of document $i$ for a given query $q$. This value function depends on the subset of documents retrieved for query $q$ (i.e., $S_{q}\subseteq D$), the summary function of documents $A(q,S)$, and the resulting performance score function $v(q,A(q,S))$. Throughout the discussion, we use the shorthand $\phi_{i}(q)=\phi_{i}(q,S_{q},A,v)$ whenever the context is clear. This setup allows the platform to aggregate document values across queries while respecting both the relative importance of documents and the frequency of the queries they support. For instance, a document that is essential for answering a rare query may be less valuable overall than a document that provides moderate value across many common queries.
We argue that it is important to conduct value attribution at the query level, especially for small publishers and niche queries. First, the economic impact of LLM-generated answers varies substantially across queries and sources. Second, large publishers can negotiate direct licensing agreements with LLM providers, but small or niche sites, whose value lies in specific, long-tail queries, cannot feasibly do so at scale. A query-level approach also ensures that marginal contributions are measured in the exact contexts where the content is used, leading to a more accurate and fair valuation. Thus, our goal is to develop a query-level document valuation approach that satisfies two properties:
@@
    Summarization Procedure Agnostic: The approach should be agnostic to the specifics of the summarization process, $A(\cdot)$, used by the platform. That is, it should generalize across RAG systems. While we present a standard RAG implementation in $\S$6.1, numerous alternatives exist-from simpler methods to more sophisticated frameworks like TextRank, GraphRAG, and DRAG (45; 14; 74). As LLM technologies rapidly evolve, our document valuation framework is designed to remain effective, regardless of advancements in summarization, ensuring broad applicability across current and future implementations.
@@
    Evaluation Process Agnostic: The approach should apply to any evaluation method, $v(\cdot)$. As discussed earlier, many explicit and implicit approaches for scoring summaries exist. Different business use cases may have access to different evaluation approaches. For example, search engines (e.g., Perplexity or Google) usually only have implicit feedback/evaluation, whereas question-answering websites or review aggregators may have more explicit feedback on the helpfulness of reviews. We would like our algorithm to be agnostic to the exact approach used. In our empirical context, we use a prompt-based LLM approach for evaluation; see $\S$6.2.
4 Solution Concept: Shapley Value for Document Valuation
This section presents a principled framework for addressing the problem of document valuation in LLM summaries, as introduced in $\S$3. In $\S$4.1, we introduce Shapley value, a game-theory concept for assigning value to individual documents in cooperative settings, and explain its relevance to our context. In $\S$4.2, we discuss the business application of using Shapley value for revenue attribution. While theoretically appealing, computing exact Shapley values is not practically applicable in most settings (including ours). To address this, $\S$4.3 introduces an efficient approximation algorithm which leverages document similarity measures to reduce the computational burden. This algorithm serves as a proof of concept within our document valuation framework; in real-world implementations, it can be adapted or extended to align with the specific architectures and operational constraints of modern LLM-based systems. Finally, in $\S$4.4, we analyze the theoretical properties of this algorithm and provide guarantees on approximation error and computation complexity.
@@ 4.1 Shapley Value
We now introduce the Shapley value along with a concise, self-contained explanation of the framework. Following the standard Shapley literature, we adapt and present the four properties of $\phi_{i}(q)$ within the context of our LLM-summarization problem to guarantee fair document valuation:
@@
    If the efficiency property is not enforced, the value function $\phi$ is only determined up to a proportional constant (see Theorem 2.1 in (20)).
@@
    Note that the combination of two queries $q_{1}+q_{2}$ does not imply that the two queries are merged into a single new query. Instead, it represents a setting where there are two distinct queries being processed (this definition naturally extends to any finite number of queries, not just two). For example, consider two queries: one on quality ($q_{1}$) and another on price ($q_{2}$). The value of document $i$ under the quality query is denoted by $\phi_{i}(q_{1})$, calculated using the performance score function $v(q_{1},A(q_{1},S))$. Similarly, $\phi_{i}(q_{2})$ represents the value of document $i$ under the price query, based on the performance score function $v(q_{2},A(q_{2},S))$. The linearity property asserts that if we sum the values obtained from these separate queries, i.e., $\phi_{i}(q_{1})+\phi_{i}(q_{2})$, the result is equivalent to the value of document $i$ calculated under a new, combined performance score function $v(q_{1},A(q_{1},S))+v(q_{2},A(q_{2},S))$. This property ensures that value attributions are invariant to the merging of multiple queries, which has important implications for revenue attribution, as discussed in §4.2.
While other desirable properties are worth discussing, these four – Efficiency, Symmetry, Null Document, and Linearity – are sufficient to uniquely determine the document value function $\phi_{i}(q)$, known as Shapley value. [^7] These four properties (axioms) are independent of any specific summarization $(A)$ or evaluation $(v)$, which are laid out in our problem definition. These properties originate from Shapley’s work on cooperative game theory (62). This is a foundational result in cooperative game theory. We refer interested readers to Shapley’s seminal paper for a formal proof, and henceforth only focus on how these properties apply to our context and their importance in our setting.
The Shapley value $\phi_{i}(q)$ for a document $i\in S_{q}\subseteq D$ can be written as the expected marginal contribution of document $i$ across all possible coalitions:
@@
Note that for $i\notin S_{q}$, the Shapley value of $i$ is $\phi_{i}(q)=0$ because only documents in $S_{q}$ are used for summarization. Documents outside $S_{q}$ have no marginal contribution and thus receive a Shapley value of zero. By simple calculation, the Shapley value formula can thus be equivalently written over the original document set $D$ as:
@@
The equivalence follows directly from the permutation-based definition, as all permutations of $D\setminus S_{q}$ do not affect the performance score.
A natural concern in attribution systems is whether document providers can strategically manipulate their content to inflate their assigned value. This issue is not new: even in traditional search engines like Google, publishers have long engaged in “mimicking the winner” strategies, which create content that resembles highly ranked pages to improve visibility. While such behavior is theoretically possible, it is difficult to execute in practice. Modern LLM-powered platforms (e.g., Perplexity, ChatGPT, Google’s AI Overview) do not disclose their specific RAG systems to external document providers. Moreover, these systems do not present explicit ranked lists but instead synthesize information into summaries, making it harder for publishers to infer their position or influence outcomes. Retrieval in such systems is governed by proprietary, black-box algorithms that incorporate a range of dynamic signals, including semantic embeddings, recency, diversity, and user engagement. This level of opacity is a deliberate design choice, observed not only in LLM systems but also in recommender platforms and digital advertising ecosystems, where hiding algorithmic details is a known strategy to deter gaming (35; 58).